# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ 5.1 MNIST
The MNIST database consists of $60000$ grayscale images of handwritten digits for the numbers 00 through $9$. Each image is of size $28\times 28$ pixels and can be viewed as a vector in $\{0,1,\dots,255\}^{784}$. Let $\mathbf{Y}$ denote the $60000\times 784$ matrix whose rows represent these images, where each row is normalized to be of unit norm. We consider a chain of $L+1$ overlapping blocks, each block corresponding to a partially observed (cosine) kernel matrix for some subset of $n=1000$ images. More specifically,
@@
Given above collection of $\{\mathbf{A}^{(i)}\}_{0\leq i\leq L}$, we compare the accuracy for jointly clustering the images in the first and last blocks. In particular, for CMMI we first construct an embedding $\hat{\mathbf{X}}^{(i)}\in\mathbb{R}^{n\times d}$ using the $d$ leading scaled eigenvectors of $\mathbf{A}^{(i)}$ for each $0\leq i\leq L$, and align $\hat{\mathbf{X}}^{(0)}$ to $\hat{\mathbf{X}}^{(L)}$ via $\breve{\mathbf{X}}^{(0)}:=\hat{\mathbf{X}}^{(0)}\mathbf{W}^{(0,1)}\cdots\mathbf{W}^{(L-1,L)}$ and $\hat{\mathbf{X}}^{(L)}$. We then concatenate the rows of $\breve{\mathbf{X}}^{(0)}$ and $\hat{\mathbf{X}}^{(L)}$ into a $2n\times d$ matrix $\mathbf{Z}^{(0,L)}$ and cluster the rows of $\mathbf{Z}^{(0,L)}$ into $3$ clusters using $K$-means. Finally we compare the accuracy of the cluster assignment against the true labels $\ell\in\{0,1,2\}$ using Adjusted Rand Index (ARI). Note that ARI values range from $-1$ to $1$, with higher values indicating closer alignment between two sets of labels. For other low-rank matrix completion algorithms, we first reconstruct $\mathbf{P}$ from $\{\mathbf{A}^{(i)}\}_{i=0}^{L}$. Letting $\hat{\mathbf{P}}$ denote the resulting estimate, we find $\hat{\mathbf{X}}$ such that $\hat{\mathbf{X}}\hat{\mathbf{X}}^{\top}$ is the best rank-$d$ approximation to $\hat{\mathbf{P}}$ in Frobenius norm (among all positive semidefinite matrices). We then subset $\hat{\mathbf{X}}$ to keep only $2n$ rows corresponding to images in $\mathbf{A}^{(0)}$ and $\mathbf{A}^{(L)}$, and finally cluster these $2n$ rows into $3$ clusters using $K$-means and compute the ARI of the resulting cluster assignments.
Comparisons between the ARIs of CMMI and other matrix completion algorithms, for different numbers of submatrices $L$, are summarized in Figure 8. Note that the black dotted line in the left panel of Figure 8 are ARIs when applying $K$-means directly on $\mathbf{Y}^{(0)}$ and $\mathbf{Y}^{(L)}$, and thus represent the best possible clustering accuracy. We observe that CMMI outperforms all competing methods on this dataset, and its ARIs remain close to optimal even as $L$ increases. Finally, the right panel of Figure 8 shows that the running time for CMMI is orders of magnitude smaller than that of other algorithms.
@@
Figure 8: The left panel reports ARIs for joint clustering of $(\hat{\mathbf{X}}^{(0)},\hat{\mathbf{X}}^{(L)})$ for subsets of the MNIST dataset using CMMI and other low-rank matrix completion algorithms. In particular, we vary $L\in\{1,2,3,4,7,9\}$ while fixing $n=1000$, $\breve{p}=0.1$, $q=0.8$, $d=2$. The black dotted line is the ARIs obtained by clustering directly on the original data (i.e, observed without noise nor missing entries). The right panel reports running time of these algorithms on a standard laptop with $8$ computing threads. The results in each panel are based on $100$ independent Monte Carlo replicates. We only evaluate the performance of GSR for $L\leq 4$ and SVT for $L=1$, as these algorithms are computationally prohibitive with even slight increases in $L$.
5.2 MEDLINE co-occurrences
The MEDLINE co-occurrences database (National Library of Medicine, 2023) summarizes the MeSH Descriptors that occur together in MEDLINE citations from the MEDLINE/PubMed Baseline over a duration of several years. A standard approach for handling this type of data is to first transform the (normalized) co-occurrence counts into pointwise mutual information (PMI), an association measure widely used in natural language processing (Church and Hanks, 1990; Lu et al., 2023). More specifically, the PMI between two concepts $x$ and $y$ is defined as $\mathrm{PMI}(x,y)=\log\tfrac{\mathbb{P}(x,y)}{\mathbb{P}(x)\mathbb{P}(y)},$ where $\mathbb{P}(x)$ and $\mathbb{P}(y)$ are the (marginal) occurrence probability of $x$ and $y$, and $\mathbb{P}(x,y)$ is the (joint) co-occurrence probability of $x$ and $y$.
For our analysis of the MEDLINE data, we first select $N=7486$ clinical concepts which were most frequently cited during the twelve years period from $2011$ to $2022$, and construct the total PMI matrix $\mathbf{P}\in\mathbb{R}^{N\times N}$ between these concepts. Next we split the dates into $L+1$ time intervals of equal length, and for each time interval $0\leq i\leq L$ we construct the individual PMI matrix $\mathbf{P}^{(i)}\in\mathbb{R}^{N\times N}$. We randomly sample, for each interval, a subset $\mathcal{U}_{i}$ of $n=1000$ concepts from those $N=7486$ cited concepts such that $|\mathcal{U}_{i-1}\cap\mathcal{U}_{i}|=100$ for all $1\leq i\leq L$. Finally we set $\mathbf{A}^{(i)}=\mathbf{P}^{(i)}_{\mathcal{U}_{i},\mathcal{U}_{i}}$ as the principal submatrix of $\mathbf{P}^{(i)}$ induced by $\mathcal{U}_{i}$. The collection $\{\mathbf{A}^{(i)}\}_{0\leq i\leq L}$ forms a chain of perturbed overlapping submatrices of $\mathbf{P}$.
Given $\{\mathbf{A}^{(i)}\}$, we apply CMMI and other low-rank matrix completion algorithms to construct $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ for the PMIs between clinical concepts in $\mathcal{U}_{0}$ and those in $\mathcal{U}_{L}$ in the total PMI matrix $\mathbf{P}$. Note that we specify $d=23$ for both CMMI and FALS, where this choice is based on applying the dimensionality selection procedure of Zhu and Ghodsi (2006) to $\mathbf{P}$. In contrast we set $d=3$ for GSR as its running time increase substantially for larger values of $d$. The values of $d$ for SVT and USVT are not specified, as both algorithms automatically determine $d$ using their respective eigenvalue thresholding procedures. We then measure the similarities between the estimated PMIs in $\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}$ and the true total PMIs in $\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}$ in terms of the Spearman’s rank correlation $\rho$ (note that, for ease of exposition, we only compare PMIs for pairs of concepts with positive co-occurrence). The Spearman’s $\rho$ between two set of vectors takes value in $[-1,1]$ with $1$ (resp. $-1$) denoting perfect monotone increasing (resp. decreasing) relationship and 00 suggesting no relationship. The results, averaged over $100$ independent Monte Carlo replicates, are summarized in Figure 9. CMMI outperforms competing algorithms in terms of both accuracy and computational efficiency.
Figure 9: The left panel reports empirical estimates of Spearman’s rank correlations for CMMI and other low-rank matrix completion algorithms as $L$ changes; we were not able to evaluate IRRSV in this experiment due to the sparse nature of the co-occurence matrix (about 60% zero entries in $\mathbf{P}$). In particular, we vary $L=\{1,2,3,5\}$ while fixing $n=1000$ and $\breve{p}=0.1.$ The right panel reports running time (in log-scale) of these algorithms on a machine with $5$ cores and $394$ GB RAM. The results in each panel are averaged over $100$ independent Monte Carlo replicates.
6 Extensions to Indefinite or Asymmetric Matrices
We now describe how the methodologies presented in this paper can be extended to block-wise data integration of symmetric indefinite matrices and asymmetric/rectangular matrices.
6.1 CMMI for symmetric indefinite matrices
Then $\mathbf{P}$ can be written as $\mathbf{P}=\mathbf{X}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}^{\top}$ with $\mathbf{X}=\mathbf{U}|\bm{\Lambda}|^{1/2}$, and thus the rows of $\mathbf{X}$ represent the latent positions for the entities. For any $i\in[K]$, let $\mathcal{U}_{i}\subseteq[N]$ denote the set of entities contained in the $i$th source, and let $\mathbf{P}^{(i)}$ be the corresponding population matrix. We then have $\mathbf{P}^{(i)}=\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}=\mathbf{U}_{\mathcal{U}_{i}}\bm{\Lambda}\mathbf{U}_{\mathcal{U}_{i}}^{\top}=\mathbf{X}_{\mathcal{U}_{i}}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}_{\mathcal{U}_{i}}^{\top}.$ For each observed submatrix $\mathbf{A}^{(i)}$ on $\mathcal{U}_{i}$, we compute the estimated latent position matrix $\hat{\mathbf{X}}^{(i)}=\hat{\mathbf{U}}^{(i)}|\hat{\bm{\Lambda}}^{(i)}|^{1/2}$. Here $\hat{\bm{\Lambda}}^{(i)}:=\mathrm{diag}(\hat{\bm{\Lambda}}_{+}^{(i)},\hat{\bm{\Lambda}}_{-}^{(i)})\in\mathbb{R}^{d\times d}$ and $\hat{\bm{\Lambda}}_{+}^{(i)}$, $\hat{\bm{\Lambda}}_{-}^{(i)}$ contain the $d_{+}$ largest positive and $d_{-}$ largest (in-magnitude) negative eigenvalues of $\mathbf{A}^{(i)}$, respectively. $\hat{\mathbf{U}}^{(i)}:=[\hat{\mathbf{U}}_{+}^{(i)},\hat{\mathbf{U}}_{-}^{(i)}]$ contains the corresponding eigenvectors.
Here $\mathcal{O}_{d_{+},d_{-}}:=\{\mathbf{O}\in\mathbb{R}^{d\times d}\mid\mathbf{O}\mathbf{I}_{d_{+},d_{-}}\mathbf{O}^{\top}=\mathbf{I}_{d_{+},d_{-}}\}$ is the indefinite orthogonal group. Eq. (6.1) implies
$\mathbf{P}_{\mathcal{U}_{1},\mathcal{U}_{2}}$ $=\mathbf{X}_{\mathcal{U}_{1}}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}_{\mathcal{U}_{2}}^{\top}=\mathbf{X}^{(1)}\mathbf{W}^{(1)}\mathbf{I}_{d_{+},d_{-}}\mathbf{W}^{(2)\top}\mathbf{X}^{(2)\top}=\mathbf{X}^{(1)}\mathbf{W}^{(1)}(\mathbf{W}^{(2)})^{-1}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}^{(2)\top}$
$=\mathbf{X}^{(1)}\mathbf{W}^{(1,2)}\mathbf{I}_{d_{+},d_{-}}\mathbf{X}^{(2)\top},$
where we define $\mathbf{W}^{(1,2)}:=\mathbf{W}^{(1)}(\mathbf{W}^{(2)})^{-1}\in\mathcal{O}_{d_{+},d_{-}}$ and we can recover $\mathbf{W}^{(1,2)}$ by aligning the latent positions for overlapping entities,
If $\mathrm{rk}(\mathbf{P}_{\mathcal{U}_{1}\cap\mathcal{U}_{2},\mathcal{U}_{1}\cap\mathcal{U}_{2}})=d$ then $\mathbf{W}^{(1,2)}=(\mathbf{X}^{(1)\top}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\mathbf{X}^{(1)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle})^{-1}\mathbf{X}^{(1)\top}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\mathbf{X}^{(2)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}=(\mathbf{X}^{(1)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle})^{\dagger}\mathbf{X}^{(2)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}$ is the unique minimizer of Eq. (6.2). Here $(\cdot)^{\dagger}$ denotes the Moore-Penrose pseudoinverse.
Now suppose $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$ are noisy observations of $\mathbf{P}^{(1)}$ and $\mathbf{P}^{(2)}$. Let $\hat{\mathbf{X}}^{(1)}$ and $\hat{\mathbf{X}}^{(2)}$ be estimates of $\mathbf{X}^{(1)}$ and $\mathbf{X}^{(2)}$ as described above. Then to align $\hat{\mathbf{X}}^{(1)}$ and $\hat{\mathbf{X}}^{(2)}$, we can consider solving the indefinite orthogonal Procrustes problem

### v4

@@ A.4 Real data experiment: MNIST
We compare the performance of CMMI against other matrix completion algorithms on the MNIST database of grayscale images. The MNIST database consists of $60000$ grayscale images of handwritten digits for the numbers 0 through $9$. Each image is of size $28\times 28$ pixels and can be viewed as a vector in $\{0,1,\dots,255\}^{784}$. Let $\mathbf{Y}$ denote the $60000\times 784$ matrix whose rows represent these images, where each row is normalized to be of unit norm. We consider a chain of $L+1$ overlapping blocks, each block corresponding to a partially observed (cosine) kernel matrix for some subset of $n=1000$ images. More specifically,
@@
Given above collection of $\{\mathbf{A}^{(i)}\}_{0\leq i\leq L}$, we compare the accuracy for jointly clustering the images in the first and last blocks. In particular, for CMMI we first construct an embedding $\widehat{\mathbf{X}}^{(i)}\in\mathbb{R}^{n\times d}$ using the $d$ leading scaled eigenvectors of $\mathbf{A}^{(i)}$ for each $0\leq i\leq L$. We specify $d=36$ for CMMI, where this choice is based on applying the dimensionality selection procedure of Zhu and Ghodsi (2006) to $\mathbf{Y}$. We then align $\widehat{\mathbf{X}}^{(0)}$ to $\widehat{\mathbf{X}}^{(L)}$ via $\breve{\mathbf{X}}^{(0)}:=\widehat{\mathbf{X}}^{(0)}\mathbf{W}^{(0,1)}\cdots\mathbf{W}^{(L-1,L)}$, and concatenate the rows of $\breve{\mathbf{X}}^{(0)}$ and $\widehat{\mathbf{X}}^{(L)}$ into a $2n\times d$ matrix $\mathbf{Z}^{(0,L)}$. We cluster the rows of $\mathbf{Z}^{(0,L)}$ into three groups using $K$-means, and evaluate clustering accuracy against the true labels $\ell\in\{0,1,2\}$ using the Adjusted Rand Index (ARI). Note that ARI values range from 1-1 to $1$, with higher values indicating closer alignment between two sets of labels. For the other low-rank matrix completion algorithms, we reconstruct $\mathbf{P}$ from $\{\mathbf{A}^{(i)}\}_{i=0}^{L}$ using $d=36$ for FALS, and $d=3$ for GSR, as the running time of GSR increases substantially with larger values of $d$. Letting $\widehat{\mathbf{P}}$ denote the resulting estimate, we then compute $\widehat{\mathbf{X}}$ such that $\widehat{\mathbf{X}}\widehat{\mathbf{X}}^{\top}$ is the best rank-$36$ approximation to $\widehat{\mathbf{P}}$ in Frobenius norm (among all positive semidefinite matrices). We extract the $2n$ rows of $\widehat{\mathbf{X}}$ corresponding to the images in $\mathbf{A}^{(0)}$ and $\mathbf{A}^{(L)}$, cluster these rows into 3 groups using $K$-means, and compute the Adjusted Rand Index (ARI) of the resulting cluster assignments.
Comparisons between the ARIs of CMMI and other matrix completion algorithms, for different numbers of submatrices $L$, are summarized in Figure A.4. We observe that CMMI outperforms all competing methods on this dataset. CMMI also has strong advantages in computational efficiency; see Section B.2 for details.
@@
Figure A.4: ARIs for joint clustering of $(\widehat{\mathbf{X}}^{(0)},\widehat{\mathbf{X}}^{(L)})$ for subsets of the MNIST dataset using CMMI and other matrix completion algorithms as we vary $L\in\{1,2,3,4\}$ while fixing $n=1000$, $\breve{p}=0.1$, $q=0.8$. The results are averaged over $100$ independent Monte Carlo replicates. We only evaluate the performance of SVT for $L\leq 2$, as these algorithms are computationally prohibitive with even slight increases in $L$.
Appendix B Integration of Multiple Matrices with Complex Connectivity
If we are given a chain of overlapping submatrices of some larger matrix $\mathbf{P}$, then Algorithm 1 provides a simple and computationally efficient procedure for recovering the unobserved regions of $\mathbf{P}$. In practice, the structure of the observed submatrices can be more complex than a simple chain. Building on the idea of Algorithm 1, which can be used to integrate any pair of connected observed submatrices, we now introduce a procedure for integrating submatrices with arbitrary overlap structures. The procedure is illustrated for positive semidefinite matrices, and it can be easily extended to the cases of symmetric indefinite matrices and asymmetric or rectangular matrices.
Suppose we have observed submatrices $\mathbf{A}^{(1)},\mathbf{A}^{(2)},\dots,\mathbf{A}^{(K)}$ for $\mathcal{U}_{1},\mathcal{U}_{2},\dots,\mathcal{U}_{K}\subset[N]$ and want to integrate them. Given Algorithm 1, a straightforward idea is to sequentially integrate each pair of connected submatrices along a chain connecting them. However, this strategy can lead to a significant amount of redundant computation. We now describe a more efficient approach that simplifies the integration process and allows all observed submatrices to be integrated simultaneously.
We consider the following undirected graph $\mathcal{G}$ to facilitate the integration process. Specifically, $\mathcal{G}$ has $K$ vertices $\{v_{1},\dots,v_{K}\}$, where each vertex $v_{i}$ corresponds to the observed submatrix $\mathbf{A}^{(i)}$ with estimated latent position matrix $\widehat{\mathbf{X}}^{(i)}$, and $v_{i}$ and $v_{j}$ are adjacent if and only if $\mathcal{U}_{i}$ and $\mathcal{U}_{j}$ are overlapping, i.e., $|\mathcal{U}_{i}\cap\mathcal{U}_{j}|\geq d$. For each pair of adjacent vertices $v_{i}$ and $v_{j}$ in $\mathcal{G}$, we can compute an orthogonal matrix $\mathbf{W}^{(i,j)}$ to align $\widehat{\mathbf{X}}^{(i)}$ and $\widehat{\mathbf{X}}^{(j)}$.
For any pair of vertices $v_{i}$ and $v_{j}$ in $\mathcal{G}$, if they are connected, meaning there exists a path between them, then the corresponding submatrices along this path form a chain that can be used to integrate $\widehat{\mathbf{X}}^{(i)}$ and $\widehat{\mathbf{X}}^{(j)}$. In the following, we assume that $\mathcal{G}$ is connected, so that all latent position estimates $\{\widehat{\mathbf{X}}^{(i)}\}_{i\in[K]}$ can be integrated. If $\mathcal{G}$ is not connected, the integration procedure can be applied separately to each connected component.
Suppose we have a graph $\mathcal{G}$ as visualized in panel (a) of Figure B.1. Note that $\mathcal{G}$ in panel (a) of Figure B.1 contains cycles, which means there exists at least one pair of vertices $v_{i}$ and $v_{j}$ with multiple paths connecting them. If, instead, there is a unique path between every pair of vertices, then all the latent position matrices $\{\widehat{\mathbf{X}}^{(i)}\}_{i\in[K]}$ can be consistently aligned, allowing all unobserved entries to be recovered simultaneously. To resolve this issue, we consider a spanning tree of $\mathcal{G}$, as illustrated in panel (b) of Figure B.1. While Theorem 2 shows that the choice of spanning tree has a negligible effect on the estimation error, we may still prefer to construct a tree such that the paths pass through vertices with smaller estimation errors. This can be achieved by setting the weight of any edge $e_{i,j}$ in $\mathcal{G}$ to $c_{i}+c_{j}$, where $c_{i}$ reflects the magnitude of the error for $\widehat{\mathbf{X}}^{(i)}$ as an estimate of $\mathbf{X}_{\mathcal{U}_{i}}$. Specifically, motivated by Lemma D.1, we define
$c_{i}:=\|(\mathbf{A}^{(i)}-\widehat{\mathbf{P}}^{(i)})\widehat{\mathbf{X}}^{(i)}(\widehat{\mathbf{X}}^{(i)\top}\widehat{\mathbf{X}}^{(i)})^{-1}\|_{F}/n_{i}^{1/2},$
where $\widehat{\mathbf{P}}^{(i)}=\widehat{\mathbf{X}}^{(i)}\widehat{\mathbf{X}}^{(i)\top}$. A minimum spanning tree (MST) of $\mathcal{G}$ is then computed based on these edge weights; see panel (c) of Figure B.1.
Figure B.1: An example of a graph $\mathcal{G}$ described in Section B (see panel (a)). Panel (b) gives a spanning tree of $\mathcal{G}$. In Panel (c), we suppose the vertices have the estimation error magnitudes $\{c_{i}\}$ as shown in the labels of vertices. The edges weights between vertices $i$ and $j$ are given by $c_{i}+c_{j}$, and we highlight the minimum spanning tree of $\mathcal{G}$ using red colored lines.
Given a minimum spanning tree, we randomly select a reference index $i_{\star}$ and align the remaining latent position estimates $\{\widehat{\mathbf{X}}^{(i)}\}_{i\neq i_{\star}}$ to $\widehat{\mathbf{X}}^{(i_{\star})}$ using the unique paths in the tree. Note that for some entities $\ell\in[N]$, we may obtain multiple estimated latent positions from different submatrices, denoted by $\{\widehat{\mathbf{X}}^{(i)}_{\langle\ell\rangle}\}_{\ell\in\mathcal{U}i}$. In such cases, we compute a weighted average of these estimates to obtain the integrated estimated latent position for entity $\ell$, where the weight assigned to each $\widehat{\mathbf{X}}^{(i)}_{\langle\ell\rangle}$ is given by $1/c_{i}^{2}$. See Algorithm B.1 for details.
Overall, the generalized CMMI procedure in Algorithm B.1 provides a principled approach for aligning entity embeddings across arbitrarily connected submatrices and for aggregating these aligned embeddings into a unified representation.
Algorithm B.1 Algorithm for holistic recovery
Step 1 Constructing the weighted graph $\mathcal{G}$:
  1. 1.
  2. 2.
    For each $i\in[K]$, obtain estimated latent positions for $\mathcal{U}_{i}$, denoted by $\widehat{\mathbf{X}}^{(i)}$, and compute $c_{i}=\|(\mathbf{A}^{(i)}-\widehat{\mathbf{P}}^{(i)})\widehat{\mathbf{X}}^{(i)}(\widehat{\mathbf{X}}^{(i)\top}\widehat{\mathbf{X}}^{(i)})^{-1}\|_{F}/n_{i}^{1/2}$.
  3. 3.
    Set the weight of each edge $e_{i,j}$ as $c_{i}+c_{j}$.
Step 2 Obtaining the aligned latent position estimates $\{\widetilde{\mathbf{X}}^{(i)}\}_{i\in[K]}$:
  1. 1.
  2. 2.